OpenStudy (anonymous):

Find the y-intercept of 2x-3y=6. Please explain. :)

OpenStudy (angela210793):

-3y=6-2x y=-2+2x/3 interception--> take x=0 and y-int=-2

OpenStudy (lalaly):

take x=0 -3y=6 y=-2

OpenStudy (saifoo.khan):

when y intercpts x = 0

OpenStudy (anonymous):

y=-2 that is (0,-2)

OpenStudy (anonymous):

Thanks! ^^

OpenStudy (saifoo.khan):

Welks! ^.^

OpenStudy (anonymous):

y=mx+b b=y intercept so you can tell from looking at 2/3 x -2 that -2 is the intercept just a tip...

OpenStudy (anonymous):

lmfao

OpenStudy (anonymous):

My mind is no longer pure.

OpenStudy (saifoo.khan):

ROFL.

OpenStudy (anonymous):

lol sorry i was gonna post this on another page but openstudy keeps getting like refresshe d everytwo seconds

OpenStudy (anonymous):

I knew it!

OpenStudy (lalaly):

haha

OpenStudy (anonymous):

<sorry toniemcm for us crashing your thread>

OpenStudy (anonymous):

I'm kind of scared to open links now. What's next, a playboy bieber? :P

OpenStudy (anonymous):

yh my bad

OpenStudy (anonymous):

It's alright. :)